# Nonlocal stress-energy tensor in time-dependent gravitational backgrounds

### v1

@@ V Energy of created particles from the RSET
When the effective action is expanded in powers of the curvature, the leading contribution to $\mathrm{Im}\,W$ is quadratic in the curvature, as shown long ago in Ref. [30] (see also Ref. [16] for a more recent discussion). However, as discussed in the previous section, the integrated flux at future null infinity obtained from the RSET vanishes. At first sight this appears puzzling, since the RSET is derived from the same effective action. We will now show that information about the energy of the created particles can nevertheless be extracted from the RSET computed to first order in the curvature. This can be done by exploiting the conservation law of the full stress–energy tensor, as described in Ref. [31].
@@
and we interpret this equation as a manifestation of energy exchange between the quantum field and the background geometry, encoded through second-order perturbative effects. Let us compute the integral of the vector $J^{\nu}$ over the entire manifold, noting that we must set $\sqrt{g}=1$ to preserve the perturbative order in $h_{\mu\nu}$. In what follows we work in Cartesian coordinates, so that ${\nabla^{\mu}}^{(0)}=\partial^{\mu}$ and ${g^{\mu\nu}}^{(0)}=\eta^{\mu\nu}$. The operator ${\nabla^{\mu}}^{(1)}$ then includes the corresponding linearized Christoffel symbols,
@@
$\int\mathrm{d}^{D-1}\vec{x}\left(\left.\braket{T^{00}}^{(2)}\right|_{t\to\infty}-\left.\braket{T^{00}}^{(2)}\right|_{t\to-\infty}\right)+\int_{\infty}^{-\infty}\mathrm{d}t\oint_{\Sigma}\mathrm{d}S_{i}\braket{T^{i0}}^{(2)}=-\int\mathrm{d}^{D}x\,J^{0}\,.$ (61)
@@
The interpretation of $\rho(p)$ as a probability density becomes more transparent when compared with the explicit computation of the particle creation probability carried out in [16], starting from the imaginary part of the effective action at second order in powers of the curvature. It was shown there that the probability of particle creation, defined as $P=1-\lvert\braket{0_{\mathrm{out}}|0_{\mathrm{in}}}\rvert^{2}$ is given by
@@
Importantly, it was also demonstrated in [16] that the probability density in Eq. (63) is strictly non-negative, ensuring that the probability of particle creation is well defined and consistent with the physical interpretation of $\Delta E$ as the corresponding energy variation. Alternatively, using the identities presented in [16], we can express Eq. (63) as
@@
To summarize, in this section we have shown that by enforcing the covariant conservation law of the second-order RSET, the total energy of the created particles can be obtained by integrating the first-order RSET contracted with local geometric quantities over the entire manifold. The final expression, Eq. (62), admits a clear physical interpretation and could, in fact, have been anticipated from previous analyses [30, 16]. Nevertheless, the derivation presented here makes explicit that this information is already encoded in the first-order RSET itself.
@@ VI Newtonian oscillating star in four dimensions
In this section, we apply our results to an example in four dimensions. We employ the Newtonian oscillating star model proposed in [16] to compute the corresponding RSET, the backreaction on the metric, and the energy exchange between the background and the quantum field. The line element reads:
@@
The Newtonian potential is assumed to be continuous at $r=a(t)$, along with its first and second $r$-derivatives to avoid singular terms in the curvature. These continuity conditions imply $f(1)=1$ and $f^{\prime}(1)=f^{\prime\prime}(1)=0$, where the primes denote derivatives with respect to the argument. For instance, a possible choice of the function $f$ satisfying the continuity constraints is an odd polynomial of fifth order: $f(x)=\frac{15}{8}x(1-\frac{2}{3}x^{2}+\frac{1}{5}x^{4})$, which has been used in Ref. [32].
@@
The leading-order contribution to the RSET can be readily computed using Eq. (41). In the limit where the duration timescale of the oscillations is much greater than their period, and this last is much greater than the radius of the star ($\tau\gg 1/\omega\gg a_{0}$), we obtain
@@
where the structure-dependent constant $\alpha_{f}$ is defined as in Ref. [16] as
@@
The leading-order term of the RSET in Eq. (72) is a Gaussian wave packet with a characteristic size of $\tau$ and a phase frequency of $\omega$. The amplitude is proportional to the total mass of the star and to the constant $\alpha_{f}$, which encodes the internal structure of the star. In contrast with time-independent scenarios [24], the leading-order term in the RSET encodes information about the internal structure of the star. The asymptotic behavior of the component $\braket{T^{t}{}_{r}}\sim 1/r^{2}$ ensures a non-vanishing energy flux through a closed surface at $r\to\infty$, and it can be readily verified that the total radiated energy vanishes when integrated over time, as expected for the RSET at first order in curvature.
@@
$h^{t}{}_{t}^{\mathrm{(q)}}=h^{t}{}_{r}^{\mathrm{(q)}}=h^{r}{}_{r}^{\mathrm{(q)}}=-\frac{4}{\pi\,r}\log\left(\mu r\right)\left[\left(\xi-\frac{1}{6}\right)^{2}+\frac{1}{180}\right]\epsilon M\omega^{4}{a_{0}}^{2}\alpha_{f}\,e^{-\frac{\pi}{2}\frac{(t-r)^{2}}{\tau^{2}}}\cos\left(\omega(t-r)\right)\,.$ (74)
@@
$\log\left(\mu r\right)\gtrsim\frac{\lambda^{4}}{a_{0}^{2}\,\,\ell_{\text{Planck}}^{2}}\,$, (75)
where $\lambda$ is the wavelength of the oscillations. This gives unrealistically large values for $r$. However, the quantum corrections for the time-dependent situation can be larger than the correction in the time independent case, which in natural units is of order $M/r^{3}$. Indeed this happens when
$r^{2}\,\log\left(\mu r\right)\gtrsim\frac{\lambda^{4}}{a_{0}^{2}}\,$, (76)
@@ VII Conclusions
We then applied this representation to compute the RSET in time-dependent geometries with spherical symmetry, concentrating on the asymptotic behavior at large distances through a multipole expansion. Our main goal was to investigate the existence of quantum hair, namely, a dependence of the RSET on the internal structure of a collapsing or oscillating star in the weak-field approximation. In previous analyses restricted to static geometries, quantum hair appears only in subleading multipoles [24, 33, 34]. In contrast, in time-dependent situations it already arises at leading order, meaning that the quantum hair manifests itself in the flux at null infinity. We obtained explicit expressions in four dimensions, showing that the RSET at null infinity can be written in terms of integrals of derivatives of the Ricci tensor over the entire past history of the source.
Although the flux at infinity is instantaneously nonvanishing, the total emitted energy vanishes. In other words, the RSET computed at first order in the curvature, and evaluated at $\mathscr{I}^{+}$, does not contain the total energy of the created particles. This fact is well known. However, the effective action from which the RSET is derived contains the information of the probability of pair creation in its imaginary part. Therefore, it should still be possible to extract the total energy of the created particles from the lowest-order RSET. We have shown explicitly that this is indeed the case, following the idea of Ref. [31]. The key point is that, although the second-order RSET is not known, it must be covariantly conserved. Expanding the conservation law in powers of the curvature reveals that the total energy is encoded in the first-order RSET. The resulting expression is closely related to the vacuum persistence probability, determined by the imaginary part of the effective action.
We also used the four-dimensional RSET to compute the backreaction on the metric. The quantum-corrected metric exhibits some unusual properties: unlike a classical gravitational wave, it does not vanish even under spherical symmetry, and its components are proportional to $\log(\mu r)\,q(t-r)/r$. Although the metric components contain the extra factor $\log(\mu r)$, we have verified that the associated stress-energy pseudo-tensor decays faster than $1/r^{2}$ as $r\to\infty$.
@@
Regarding future work, the methods and results presented here could be extended to charged and/or slowly rotating configurations, still within the weak-field approximation. Beyond this regime, it would be interesting to compare our findings with those obtained from other nonlocal approaches, such as spherical dimensional reduction with the subsequent use of the $1+1$ Polyakov effective action.
@@ Acknowledgments
This research was supported by Consejo Nacional de Investigaciones Científicas y Técnicas (CONICET) and Universidad Nacional de Cuyo (UNCUYO).

### v2

@@ V Energy of created particles from the RSET
When the effective action is expanded in powers of the curvature, the leading contribution to $\mathrm{Im}\,W$ is quadratic in the curvature, as shown long ago in Ref. [Frieman1985] (see also Ref. [nosotros_cotton] for a more recent discussion). However, as discussed in the previous section, the integrated flux at future null infinity obtained from the RSET vanishes. At first sight this appears puzzling, since the RSET is derived from the same effective action. We will now show that information about the energy of the created particles can nevertheless be extracted from the RSET computed to first order in the curvature. This can be done by exploiting the conservation law of the full stress–energy tensor, as described in Ref. [horowitz].
@@
and we interpret this equation as a manifestation of energy exchange between the quantum field and the background geometry, encoded through second-order perturbative effects. Let us compute the integral of the vector $J^{\nu}$ over the entire manifold, noting that we must set $\sqrt{g}=1$ to preserve the perturbative order in $h_{\mu\nu}$. In what follows we work in Cartesian coordinates, so that $\nabla_{\mu}^{(0)}=\partial_{\mu}$ and $g_{\mu\nu}^{(0)}=\eta_{\mu\nu}$. The operator $\nabla_{\mu}^{(1)}$ then includes the corresponding linearized Christoffel symbols,
@@
$\int\mathrm{d}^{D-1}\vec{x}\left(\left.\braket{T^{00}}^{(2)}\right|_{t\to\infty}-\left.\braket{T^{00}}^{(2)}\right|_{t\to-\infty}\right)+\int_{-\infty}^{\infty}\mathrm{d}t\oint_{\Sigma}\mathrm{d}S_{i}\braket{T^{i0}}^{(2)}=-\int\mathrm{d}^{D}x\,J^{0}\,.$ (59)
@@
The interpretation of $\rho(p)$ as a probability density becomes more transparent when compared with the explicit computation of the particle creation probability carried out in [nosotros_cotton], starting from the imaginary part of the effective action at second order in powers of the curvature. It was shown there that the probability of particle creation, defined as $P=1-\lvert\braket{0_{\mathrm{out}}|0_{\mathrm{in}}}\rvert^{2}$ is given by
@@
Importantly, it was also demonstrated in [nosotros_cotton] that the probability density in Eq. (61) is strictly non-negative, ensuring that the probability of particle creation is well defined and consistent with the physical interpretation of $\Delta E$ as the corresponding energy variation. Alternatively, using the identities presented in [nosotros_cotton], we can express Eq. (61) as
@@
To summarize, in this section we have shown that by enforcing the covariant conservation law of the second-order RSET, the total energy of the created particles can be obtained by integrating the first-order RSET contracted with local geometric quantities over the entire manifold. The final expression, Eq. (60), admits a clear physical interpretation and could, in fact, have been anticipated from previous analyses [Frieman1985, nosotros_cotton]. Nevertheless, the derivation presented here makes explicit that this information is already encoded in the first-order RSET itself.
@@ VI Newtonian oscillating star in four dimensions
In this section, we apply our results to an example in four dimensions. We employ the Newtonian oscillating star model proposed in [nosotros_cotton] to compute the corresponding RSET, the backreaction on the metric, and the energy exchange between the background and the quantum field. The line element reads:
@@
The Newtonian potential is assumed to be continuous at $r=a(t)$, along with its first and second $r$-derivatives to avoid singular terms in the curvature. These continuity conditions imply $f(1)=1$ and $f^{\prime}(1)=f^{\prime\prime}(1)=0$, where the primes denote derivatives with respect to the argument. For instance, a possible choice of the function $f$ satisfying the continuity constraints is an odd polynomial of fifth order: $f(x)=\frac{15}{8}x(1-\frac{2}{3}x^{2}+\frac{1}{5}x^{4})$, which has been used in Ref. [estrella_gas].
@@
The leading-order contribution to the RSET can be readily computed using Eq. (41). In the limit where the duration timescale of the oscillations is much greater than their period, and this latter is much greater than the radius of the star ($\tau\gg 1/\omega\gg a_{0}$), we obtain
@@
where the structure-dependent constant $\alpha_{f}$ is defined as in Ref. [nosotros_cotton] as
@@
The leading-order term of the RSET in Eq. (70) is a Gaussian wave packet with a characteristic size of $\tau$ and a phase frequency of $\omega$. The amplitude is proportional to the total mass of the star and to the constant $\alpha_{f}$, which encodes the internal structure of the star. In contrast with time-independent scenarios [fdm_vacuum_polarization], the leading-order term in the RSET encodes information about the internal structure of the star. The asymptotic behavior of the component $\braket{T^{t}{}_{r}}\sim 1/r^{2}$ ensures a non-vanishing energy flux through a closed surface at $r\to\infty$, and it can be readily verified that the total radiated energy vanishes when integrated over time, as expected for the RSET at first order in curvature.
@@
$h^{t}{}_{t}^{\mathrm{(q)}}=h^{t}{}_{r}^{\mathrm{(q)}}=h^{r}{}_{r}^{\mathrm{(q)}}=\frac{4}{\pi\,r}\left[\left(\xi-\frac{1}{6}\right)^{2}+\frac{1}{180}\right]\epsilon M\omega^{4}{a_{0}}^{2}\alpha_{f}\,e^{-\frac{\pi}{2}\frac{(t-r)^{2}}{\tau^{2}}}\cos\left(\omega(t-r)\right)\,.$ (72)
@@
$\ell_{\text{Planck}}\gtrsim\frac{\lambda^{2}}{a_{0}}\,$, (73)
where $\lambda$ is the wavelength of the oscillations. This gives unrealistically small values for the ratio $\lambda^{2}/a_{0}$. However, the quantum corrections for the time-dependent situation can be larger than the correction in the time independent case, which in natural units is of order $M/r^{3}$. Indeed this happens when
$r\gtrsim\frac{\lambda^{2}}{a_{0}}\,$, (74)
@@ VII Conclusions
We then applied this representation to compute the RSET in time-dependent geometries with spherical symmetry, concentrating on the asymptotic behavior at large distances through a multipole expansion. Our main goal was to investigate the existence of quantum hair, namely, a dependence of the RSET on the internal structure of a collapsing or oscillating star in the weak-field approximation. In previous analyses restricted to static geometries, quantum hair appears only in subleading multipoles [fdm_vacuum_polarization, Calmet2019, Calmet2021]. In contrast, in time-dependent situations it already arises at leading order, meaning that the quantum hair manifests itself in the flux at null infinity. We obtained explicit expressions in four dimensions, showing that the RSET at null infinity can be written in terms of integrals of derivatives of the Ricci tensor over the entire past history of the source.
Although the flux at infinity is instantaneously nonvanishing, the total emitted energy vanishes. In other words, the RSET computed at first order in the curvature, and evaluated at $\mathscr{I}^{+}$, does not contain the total energy of the created particles. This fact is well known. However, the effective action from which the RSET is derived contains the information of the probability of pair creation in its imaginary part. Therefore, it should still be possible to extract the total energy of the created particles from the lowest-order RSET. We have shown explicitly that this is indeed the case, following the idea of Ref. [horowitz]. The key point is that, although the second-order RSET is not known, it must be covariantly conserved. Expanding the conservation law in powers of the curvature reveals that the total energy is encoded in the first-order RSET. The resulting expression is closely related to the vacuum persistence probability, determined by the imaginary part of the effective action.
We also used the four-dimensional RSET to compute the backreaction on the metric. The quantum-corrected metric exhibits some interesting properties: unlike a classical gravitational wave, it does not vanish even under spherical symmetry. Moreover, although the metric components decay as $1/r$, we have verified that the associated stress-energy pseudo-tensor decays faster than $1/r^{2}$ as $r\to\infty$.
@@
Regarding future work, the methods and results presented here could be extended to charged and/or slowly rotating configurations, still within the weak-field approximation. Beyond this regime, a possible avenue is to explore generalizations of the nonlocal effective action applicable to strong gravitational fields. Since this is a technically demanding task (the third-order corrections already display considerable complexity [Barvinsky:1994]) several approaches have been developed to partially incorporate the quantum effects. For example, long time ago, Vilkovisky and collaborators proposed an effective action for conformal fields in which the nonlocal form factors $\beta(\Box)$ are evaluated on a conformally flat metric rather than on flat spacetime [Fradkin:1978, Vilkovisky:1984]. A local approximation to this effective action has been employed in analyses of gravitational collapse [Frolov:1981]. In a similar direction, various authors have considered nonlinear completions of the second-order effective action, where the nonlocal form factors are evaluated in cosmological backgrounds [Hamber:2005, lopez_nacir_fdm, Donoghue:2014]. Other approaches to go beyond the weak-field regime include performing a spherical dimensional reduction and subsequently applying the $(1+1)$-dimensional Polyakov effective action [Fabbri:2005], as well as exploring its generalization to $3+1$ dimensions through the Riegert action [Riegert:1984, Mottola:2006, Shapiro:2008sf]. A systematic comparison and possible extensions of these frameworks are currently underway.
@@ Acknowledgments
This research was supported by Consejo Nacional de Investigaciones Científicas y Técnicas (CONICET).